Let S ⊆ \Re be the tetrahedron having vertices (0, 0, 0), (0, 1, 1), (1, 2, 3), and (−1, 0, 1).

Let f : \Re ^{3}\Re ^{} be the function defined by f(x, y, x) = x − 2y + 3z.

Using the change of variables theorem,

rewrite \int_s f as an integral over a 3-rectangle, then use Fubini’s theorem to evaluate the integral.
Answer #1

ANSWER:-ô C -25 [ the given information is S; of (2, 4, Z) = x-2y+32 change of variables x = 5, y =5, zat - d (hy,z) dir, sit) dk dor* །-7 a N D + 》། 《-༽ཀུ + སྐྱེ+4 AS - + s + རྟེ)as 47 ) [= - (3) [ + ༥ |-དུ་ ༈ : , , [༌ ཀ -27/1 - ) + ་ [t - ] + a7 || ལ) – (